OpenStudy (anonymous):

10 • 4n – 1 Find the first four terms of the multiplication pattern given by the formula.

OpenStudy (azureilai):

You substitute I'm assuming your first term is 1 10*4(1)-1= 10*4(2)-1= 10*4(3)-1= 10*4(4)-1=

OpenStudy (anonymous):

i dont get it :/